OpenStudy (anonymous):

\[\sqrt{4x+5}-\sqrt{x-1}=3\] please help need step by step instruct

OpenStudy (curry):

\[\sqrt{4x+5}/\sqrt{x-1}=3\] is that the right equation

OpenStudy (anonymous):

yes

OpenStudy (curry):

clear the radical \[\sqrt{4x^2-x-5}/x-1 = 3\]

OpenStudy (curry):

\[2x \sqrt{-x-5}=3x-1\]

OpenStudy (curry):

ill let you solve it from there

OpenStudy (anonymous):

how did you get 4x^s-x-5

OpenStudy (curry):

cause i multiplied 4x-5 by x-1

OpenStudy (anonymous):

oook i can see that there r 2 possible answers one of them being 1 and the aquare root of 4 is 2 so do i use the quadratic formula or aquare root method

OpenStudy (radar):

is it \[\sqrt{4x+5}/\sqrt{x-1}=3\] Or is it\[\sqrt{4x+5}-\sqrt{x-1}=3\] These are two very different kinds of problem.

OpenStudy (radar):

If it is the first that is pretty easy. square both sides getting (4x+5)/(x-1)=9 cross multiply getting: 9x-9=4x+5 5x=14 x=14/5

OpenStudy (radar):

if it the second then I need to think about it lol

OpenStudy (radar):

O.K if it is the subtraction (the second one) here is the solution: x+5 and x=1\[\sqrt{4*5 +5}-\sqrt{5-1}=\sqrt{25}-\sqrt{4}=5-2=3\] \[\sqrt{4*1+5}-\sqrt{1-1}=\sqrt{9}-\sqrt{0}=3-0=3\]

OpenStudy (radar):

If you want to see the steps (kind of wobbly) to obtain the solution, let me know below.

OpenStudy (anonymous):

its the second one and the answer is one and 5 but how did they get the answer

OpenStudy (radar):

O.K. I am back. Here are the steps that I took to solve this thing. There may be other ways to solve this. I am having connection problems. Can't even get the equation bar to work. will try later

OpenStudy (radar):

Here goes w/o editor. I will try and write the equation again so we have it all in one spot. (4x+5)^1/2 -(x-1)^1/2 = 3 sq rt (4x+5) - sq rt (x-1) = 3 Step 1. (4x+5)^1/2 = 3 + (x-1)^1/2 Adding sq rt (x-1) to both sides. Step 2. Square both sides getting: 4x+5=9+6(x-1)^1/2 + x - 1 now combine and transpose to get 3x-3 = 6(x-1)^1/2 Divide thru by 3 getting x-1=2(x-1)^1/2 Step 3. Square both sides getting x^2-2x+1=4(x-1) combine and simplify getting x^2-6x+5=0 Step 4. Factor getting: (x-5)(x-1) = 0 or x-5=0 giving x=5 x-1=0 giving x=1 Hope this helps.